my question is this:
There were 44,000 people at a ball game in Los Angeles. The day's receipts were $369,000. How many people paid $13.00 for reserved seats and how many paid $6.00 for general admission?
in the class, we are working with using systems of equations to combine equations and eliminate one variable to figure out the other variable ie:
x + y = 5
x - y = -9
2x = -4
x = -2
and those types of problems. for this question, i can not figure out what to use as variables for either side of the equation. the other word question we were given gave 6 variables, so we were able to plug in a numerical value to each section. here- theres only 4 numbers.
is this a different type of question entirely, or- am i just missing the point entirely?
thanks in advance!

Answer by belle3833(46)

x be the number of people that paid$13 and y be the number of people that paid $6
so x+y=44000 (the total number of people that attended)
13x+6y=369000 (total revenue)
These are your two equations. YOu need to solve them simulateously. If you have problems completing it from here please message me